When Sarah turns 36 years old, her father will be 62. Currently, Sarah is 1/3 of her father's age. How many years old is Sarah now?

Found 2 solutions by macston, stanbon:
Answer by macston(5194)

S=Sarah's age; F=Father's age=S+26yrs
.
F=3S
S+26yrs=3S
26yrs=2S
13yrs=S ANSWER: Sarah is currently 13 years old.

Answer by stanbon(75887)

When Sarah turns 36 years old, her father will be 62. Currently, Sarah is 1/3 of her father's age. How many years old is Sarah now?
--------
Equations:
s = (1/3)f
----
Father is always 62-36 = 26 years older than Sarah.
------
f = s + 26
-----
Substitute for "f" and solve for "s"::
s = (1/3)(s+26)
3s = s+26
2s = 26
s = 13 yrs (Sarah's age now)
-------------------
Cheers,
Stan H.
----------------